OpenStudy (anonymous):

PLEASE !!! write an equation for a line with a slope m= -2/7 and a Y intersept at (0,0)

OpenStudy (anonymous):

y-0=-2/7(x-0) y=(-2/7)x

OpenStudy (anonymous):

oh thank you

OpenStudy (anonymous):

is that the whole process?

OpenStudy (anonymous):

yup thats it